OpenStudy (anonymous):

The first angle of a triangle is 4° larger than twice the second angle. The third angle is 21° smaller than twice the first angle. Find the measure of the three angles. Write your answers in the order of 1st angle, 2nd angle, 3rd angle.

OpenStudy (anonymous):

There are three unknowns. lets called them A, B, C. They are first, second and third angle. For three unknown we need three equations: First angle, A equals: A = 2B +R Third angle, C: C = 2A - 21 The sum of all angles: A + B + C = 180. Solve the three equations.

OpenStudy (radar):

Let x=Second angle then 2x+4 = First angle (given information) then 2(2x+4)-21 = 3rd angle. What do we know about the sum of angles in a triangle? The sum is 180 degrees then x +2x+4 +2(2x+4)-21 =180 simplifying 7x=189 or x=27 degrees Second angle is 27 degrees. Leaving the first and third angle for you as practice makes perfect!

OpenStudy (anonymous):

Sorry, it is A= 2B + 4

OpenStudy (anonymous):

u did well tenxz,what is A=2B+ 4 fits in the problem